OpenStudy (aroub):

The length of the diagonal of a rectangle with length x and width x - 1 is:

OpenStudy (anonymous):

Can we use Pythagorean theorem again?

OpenStudy (aroub):

Yes, it's all related to Pythagorean Theorem

OpenStudy (anonymous):

Do you know that the each internal angles of a rectangle is \(90^\circ \)

OpenStudy (mertsj):

\[\sqrt{x^2+(x-1)^2}=\sqrt{x^2+x^2-2x+1}=\sqrt{2x^2-2x+1}\]

OpenStudy (aroub):

YES!

OpenStudy (anonymous):

So, can you solve it now arouba?

OpenStudy (aroub):

yes :

OpenStudy (aroub):

:)

OpenStudy (anonymous):

Aroub is smart :)

OpenStudy (aroub):

Obviously :P